OpenStudy (anonymous):

how do i make 2.25 into an improper fraction?

OpenStudy (amorfide):

you have a whole number, 2 2 is 2x1 1/1 is a fraction for 1 so you have 2 of these giving you 2/1=2 you have 0.25 left over you want a fraction that gives 0.25 0.25x=1 x=1/0.25 x=1/4 you have 2 + 1/4 =2.25 now you have \[\frac{ 2 }{ 1 } + \frac{ 1 }{ 4 }\] just add fractions

OpenStudy (anonymous):

Thank you so much I am not that good at math

OpenStudy (amorfide):

to add fractions \[\frac{ 2 }{ 1 } + \frac{ 1 }{ 4 } = \frac{ (2 \times 4)+(1 \times 1) }{ 1 \times 4 }\] because you cross multiply

OpenStudy (anonymous):

2.25 in a fraction would be 2 and 1/4 or it could be 9/4

OpenStudy (amorfide):

9/4 which is what i gave her, that is improper form 2 and 1/4 is mixed form

OpenStudy (anonymous):

ik

OpenStudy (anonymous):

thnx!

OpenStudy (anonymous):

gave u one 2 :)

OpenStudy (amorfide):

ty , lack of medals lately people just cba

OpenStudy (anonymous):

ikr! aint got one in 4 days

OpenStudy (anonymous):

im yur fan